Bill is 1 year less than twice as old as his brother John. If the sum of their ages is 20 years, how old are Bill and his brother John. How do I solve this?
Answer by Cromlix(4381)

Bill is 1 year less than twice his brother, John.
B = (2J) - 1
B + J = 20
Replace B with (2J) - 1
2J - 1 + J = 20
3J = 20 + 1
3J = 21
J = 7
John = 7 years old
Bill = 13 years old
Hope this helps.
:-)